I have a question that states: evaluate each the function and simplify the results for each of the following:
f(x)={1/2x if x<-3
{1-3x if x > or equal to -3
simplify:
a)f(-3)
b)f(0)
c)f(4)
d)f(-2.1)
PLease help, I have no clue how to do these 4 problems....

Answer by stanbon(48535)

f(x)={1/2x if x<-3
{1-3x if x > or equal to -3
Comment: Notice you have two rules for determining f(x); which rule you
use depends of the value of x.
---------------------------------------
simplify:
a)f(-3)
Since -3 is >=-3 you want to use f(x)=1-3x; So, f(-3)=1-3(-3)=10
----------------
b)f(0)
Since 0 is >=-3 you want to use f(x)=1-3x; So, f(0)=1-3*0 = 1
------------------
c)f(4)
Since 4 is >=-3 you want to use f(x)=1-3x; So f(4)=1-3*4=-11
----------------
d)f(-2.1)
Since -2.1 is >=-3 you want to use f(x)=1-3x; So f(2.1)=1-3(2.1)= 1-6.3=-5.3
=============
Cheers,
Stan H.